# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ III-E5 Estimation Variants of CFR
After advantage-based regret minimization (ARM) [88] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR [89] utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR [90] leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and its proven effectiveness in larger-scale poker games. Moreover, Single Deep CFR (SD-CFR) [91] demonstrates that training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM [92] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) [93] combines the thought of retrospective policy improvement with CFR.
III-F Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section and illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general. We also summarize the key points in Table II.
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and as training progresses, the policy population expands. Therefore, in our framework, we use placeholder initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes, designed to handle more intricate tasks. For example, $\alpha$-PSRO [46] specifically utilizes an $\alpha$-rank-based MSS to address multi-player general-sum games.
Unlike the two previously mentioned categories, the ongoing-training-based series employs a different paradigm by training the entire policy population together. This approach aims to strengthen all policies simultaneously at each epoch, rather than expanding the policy population and relying on newer policies to be stronger. As a result, actual initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential. Consequently, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
Lastly, the regret-minimization-based series focuses on the overall performance over time rather than on a single episode. For example, it is well-suited for games like Texas Hold’em, which require strategies involving deception, concealment, and bluffing. The main goal of the training process in this series is to update the regrets associated with different strategies. In our framework, we use $h(i)$ to store this information. Since the policies are determined by $h(i)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4). We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
Our framework is built upon PSRO [42] and NeuPL [43]. Here, we outline the differences between our framework and these two existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$, rather than being the meta-strategy of policy $i$. Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 4 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
@@ III-F Reassessment of the Framework
Table II: Overview of Main Self-play Algorithms within Our Framework.
Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$
@@
Vanilla Self-play [6] Equ. (33) $\emptyset$ Traditional Self-play $E=1$ Placeholder General
@@
Independent RL Equ. (37)
DO [45] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Placeholder General
@@
FTW [70] Equ. (38) $\emptyset$ Ongoing-training-based $E>1$ Actual $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$
@@
Vanilla CFR [72] Equ. (33) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Placeholder $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$
@@
IV Empirical Analysis
@@
In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games and Mahjong, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play. We then illustrate how self-play is applied in each of these complex scenarios and provide a comparative analysis of these applications in Table III.
IV-A Board Games
@@
The landscape of board games, the majority of which are perfect information games, was previously revolutionized by the introduction of two key techniques: position evaluation and Monte Carlo tree search (MCTS) [94, 95]. These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [96], checkers [97], othello [98], backgammon [99], and Scrabble [100]. In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level [101, 102, 103, 104, 105]. In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play. In addition to Go, we will broaden our exploration to include Stratego, a board game characterized by imperfect information, contrasting the majority of board games that are based on perfect information.
IV-A1 Go
The paradigm of Go is revolutionized with the launch of DeepMind’s AlphaGo series [8, 9, 10, 11], which leveraged the power of self-play to significantly elevate performance, setting a new benchmark in the field of Go.
@@
In AlphaGo [8], the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS expansion and a precise policy network $p_{\sigma}(a|s)$. The second stage employs RL to refine the policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Self-play is instrumental here. More specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly chosen historical version $p_{\rho^{-}}(a|s)$, akin to the MSS shown in Equ. (34). Afterwards, $v_{\theta}(s)$ is trained using the samples from the games in which the trained policy network $p_{\rho}(a|s)$ competes against itself. In the final stage, MCTS integrates the policy and value networks to inform action selection.
@@
Unlike AlphaGo, AlphaGo Zero [9] does not require any expert data except game rules and instead learns through self-play. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $p=\text{Pr}(a|s)$ and the state value $v$. Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, as opposed to the rollouts used in AlphaGo. Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (37). Furthermore, for a new policy to be incorporated into the policy pool, it must surpass a 55 percent win rate against its predecessor, aligning with the stipulations set in Algo. 2 at Line 1.
AlphaZero [10] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi, with some modifications. Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi. Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero.
Building upon AlphaZero, MuZero [11] takes the concept of learning from scratch to the next level, even operating without predefined game rules. MuZero incorporates ideas from model-based RL to model the dynamics of games. More concretely, in addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states. These three networks are trained jointly. Similarly to AlphaGo Zero and AlphaZero, MuZero employs MCTS guided by the three aforementioned networks to make decisions. The self-play process in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games. Furthermore, several studies have extended the capabilities of MuZero. For example, Stochastic MuZero [106] learns a stochastic model instead of a deterministic model to enhance its performance in more complex scenarios. Sampled MuZero [107] makes learning feasible in games with intricate action spaces.
IV-A2 Stratego
@@
Unlike most board games which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and a large number of potential game states, estimated $10^{535}$ [108]. The game is divided into two phases: the deployment phase, where players secretly arrange their units, setting the stage for strategic depth, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [108] showed promising advances in AI’s ability to tackle it.
DeepNash scales up R-NaD [54] to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. In the dynamics stage, Neural Replicator Dynamics [109] is utilized to obtain the approximate fixed-point policy effectively. Additionally, DeepNash incorporates train-time fine-tuning and test-time post-processing methods to eliminate unreliable actions, thereby enhancing the robustness and accuracy of its output. DeepNash holds the third-place ranking among all Gravon Stratego players and secures victories in almost every match against existing Stratego bots.
IV-B Card Games and Mahjong
IV-B1 Texas Hold’em
@@
Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements. In its two-player variant, it’s referred to as heads-up Texas Hold’em. The game becomes more complex with additional players. The gameplay begins with each player receiving two private cards (hole cards), followed by a round of betting. Subsequently, three community cards (the flop) are revealed, leading to another betting round. This is followed by the dealing of a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds. The objective is to construct the best five-card poker hand from any combination of hole cards and community cards. The betting proceeds until a showdown occurs, where the remaining players disclose their cards. The individual with the superior hand claims the pot. Moreover, Texas Hold’em is distinguished by its three betting formats: no-limit, fixed-limit, and pot-limit. Each format offers a different risk profile and complexity in terms of game state sizes. Among these, the no-limit format is particularly noted for its complexity, allowing players to bet any amount up to their entire stack of chips. Its popularity is partly due to its frequent appearance in tournaments and media, making it a staple in modern poker culture. While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis and educational contexts within game theory, our focus will focus on the development of algorithms designed to compete in real-world Texas Hold’em.
In poker, especially Texas Hold’em, abstraction is crucial to simplifying the inherent complexity of the game [110]. This methodology reduces the vast state space of the game to a manageable size, thus allowing algorithms to calculate strategies more efficiently. There are two main types of abstraction in Texas Hold’em: action abstraction, which simplifies the range of possible moves, and information abstraction, which groups similar hand strengths or pot sizes, streamlining the decision-making process.
@@
Although heads-up limit Texas Hold’em (HULHE), the simplest form of Texas Hold’em with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [111]. It utilizes fixed-point arithmetic with compression and CFR+ [75] to address the issues of storage and computation, respectively, thus solving HULHE without involving any neural networks. This resulted in superhuman performance and surpassed previous agents. Moreover, NSFP [56] introduces the use of end-to-end RL through self-play training, achieving competitive performance in HULHE. Similarly, Poker-CNN [112], which utilizes a 3D-vector representation of HULHE combined with convolutional networks, learned through self-play and also achieved competitive performance. Deep CFR [90], combines deep neural networks with CFR to surpass the performance of NFSP. However, it’s noteworthy that despite advancements, network-based methods haven’t surpassed the regret-based method used by Cepheus, despite its substantial computational demands.
After the solution of HULHE, research focus shifted to heads-up no-limit Texas Hold’em (HUNL), representing a more complex challenge due to its significantly larger space of game states, approximately $10^{164}$ [113]. Thus, it is impossible to traverse the game tree as Cepheus does in HULHE. The primary technique employed by DeepStack [12] is continual re-solving. This method involves dynamically computing the strategy in real-time when it is the program’s turn to act, focusing on a subtree of limited depth and breadth. To estimate the outcomes of the furthest reaches of this subtree, DeepStack utilizes neural networks. These networks take into account both the player’s range and the opponent’s counterfactual values, which are monitored throughout the entire game. The neural networks employed include the turn network, the flop network, and the auxiliary network, all of which are pre-trained using 10 million randomly generated poker games. The target counterfactual values for these generated poker situations are derived through thousands of iterations using CFR+ [75]. DeepStack has demonstrated competitive performance against professional poker players by integrating these approaches. Libratus [86] employs self-play to develop its blueprint strategy based on an abstraction of the game, leveraging an enhanced version of MCCFR [77]. Initially, the blueprint strategy is deployed for decision-making in the first two rounds, benefiting from the detailed abstraction. In the later rounds, the blueprint strategy aids in nested safe subgame solving, akin to DeepStack’s continual re-solving aiming to get a more precise strategy for subgames in real-time. Additionally, Libratus features a self-improvement component designed to enhance the blueprint strategy by addressing any missing branches. Through these strategies, Libratus has proven capable of defeating professional HUNL players. ReBel [114] expands the concept of a game state in poker into a public belief state (PBS), thereby transforming the imperfect-information game into a perfect-information game with continuous state space. ReBel adopts the methodology of AlphaZero [10], which combines RL with a search algorithm to train both the value network and the policy network by self-play. To mitigate heavy computation and storage demands, AlphaHoldem [115] employs an end-to-end self-play RL method augmented with several additional techniques, to outperform its reimplementation of DeepStack. During training, the system employs the $K$-best Self-play, where it selects the top $K$ agents based on their ELO ratings [35] and the newly initialized agent. It then generates experience replays from competitions among these $K+1$ agents.
@@
When there are more than two players in Texas Hold’em, the game becomes significantly more complex. Pluribus [116] based on the framework of Libratus [86], addresses this complexity in six-player no-limit Texas Hold’em. It employs both action abstraction and information abstraction to simplify the intricate dynamics of the original game. Similarly to Libratus, Pluribus utilizes MCCFR [77] in self-play to develop its blueprint strategy. Building on this blueprint strategy, it conducts a depth-limited search before executing actions. More specifically, diverging from Libratus, Pluribus maintains a streamlined policy pool consisting of only four strategies. This approach assumes that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently. Despite its strategy not adapting to the observed tendencies of its opponents, and lacking theoretical proofs of achieving an NE in six-player scenarios, Pluribus has demonstrated its capability by outperforming top professional players in six-player no-limit Texas Hold’em.
IV-B2 DouDiZhu

### v2

@@ 3.5.5 Estimation Variants of CFR
Although these CFR variants advance the field, they can’t directly solve large imperfect-information extensive-form games due to their reliance on tabular representations. The typical approach involves abstracting the original game, applying CFR to the abstracted game, and translating strategies back to the original. This abstraction is game-specific and relies heavily on domain knowledge. Additionally, smaller abstractions often yield suboptimal results. Given these challenges, [133] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, using regression trees as the regressor limits RCFR’s applicability to miniature games, and the necessity for manually crafted features remains a drawback. After advantage-based regret minimization (ARM) [62] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR [75] utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR [14] leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games. Single Deep CFR (SD-CFR) [118] demonstrates training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM [119] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) [46] combines the thought of retrospective policy improvement with CFR.
@@
Table 1: Overview of Representative Self-play Algorithms within Our Framework.
Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$ Suitable Game
@@
Vanilla Self-play [108] Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
@@
DO [92] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
@@
FTW [60] Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
@@
Vanilla CFR [152] Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
@@
3.6 Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section, explain the rationale behind categorizing the algorithms, and summarize the representative algorithms in Table 1. Moreover, we will illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general.
3.6.1 Categorization Rationale
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and the policy population expands as training progresses. Therefore, in our framework, we use lazy initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes to handle tasks with intricate game-theoretical requirements. For example, $\alpha$-PSRO [96] specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games. In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSes in the PSRO series.
Unlike the two previously mentioned categories, the ongoing-training-based series adopts a different paradigm. Instead of gradually expanding the policy population and relying on newer policies to be stronger, this approach strengthens all policies simultaneously at each epoch. This method alleviates issues such as early truncation and repeated skill relearning that occur in the above two categories. To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential. Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
Lastly, the regret-minimization-based series focuses on overall performance over time rather than on individual episodes, making it particularly suitable for repeated games. For example, Texas Hold’em is a classic repeated game where players adjust their strategies based on past interactions and use tactics involving deception and bluffing. The main goal of this training process is to update the regrets associated with different strategies, which in turn demands significant storage resources. Our framework uses $h(i)$ to store this information. Since the policies are determined by $h(i)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4). We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
3.6.2 Comparing with Existing Frameworks
Our framework is built upon PSRO [73] and NeuPL [82]. Here, we outline the differences between our framework and these existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$ rather than being the meta-strategy of policy $i$. Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 5 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
4 Empirical Analysis
@@
In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play. We then illustrate how self-play is applied in these complex scenarios and provide a comparative analysis of these applications in Table 2.
4.1 Board Games
@@
Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) [29, 67]. These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [24], checkers [110], othello [23], backgammon [126], and Scrabble [114]. In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level [9, 30, 6, 37, 43]. In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play. In addition to Go, we will explore Stratego, a board game with imperfect information, in contrast to most board games that only involve perfect information.
4.1.1 Go
@@
Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory. The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series [115, 117, 116, 112], which leveraged the power of self-play to significantly elevate performance. In AlphaGo [115], the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$. The second stage employs self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly selected historical version $p_{\rho^{-}}(a|s)$, similar to the MSS shown in Equ. (6), while $v_{\theta}(s)$ is trained using the game samples collected by self-play of $p_{\rho}(a|s)$. In the third stage, MCTS integrates the policy and value networks to select actions.
@@
Based on AlphaGo, AlphaGo Zero [117] does not require any expert data except game rules. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities and the state value. Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (5). A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor. AlphaZero [116] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi. Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network to generate samples by self-play without validation. Building upon AlphaZero, MuZero [112] takes learning from scratch to the next level, even operating without predefined game rules. The self-play process in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games, MuZero also achieves state-of-the-art performance in Atari games.
4.1.2 Stratego
@@
Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ [104]. The game is divided into two phases: the deployment phase, where players secretly arrange their units, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [104] showed promising advances in AI’s ability to tackle it. DeepNash scales up evolution-theory-based self-play method R-NaD [105] (discussed in Sec. 3.3.5) to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. Neural Replicator Dynamics [52] is utilized to obtain the approximate fixed-point policy. DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
4.2 Card Games
Unlike board games, card games are typically imperfect information games with a larger state space and greater complexity. Here, we introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games. Texas Hold’em is a repeated game rather than a stage game, so the focus is on the overall win-loss outcome rather than on individual rounds. Consequently, it is particularly well-suited for regret-minimization-based self-play methods, which aim to minimize overall regret rather than regret on a per-round basis. In contrast, DouDiZhu and Mahjong are typically regarded as stage games, so studies often employ self-play methods that do not rely on regret minimization.
4.2.1 Texas Hold’em
@@
Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements. The two-player variant is called heads-up Texas Hold’em. The gameplay begins with each player receiving two private cards (hole cards), followed by a round of betting. Subsequently, three community cards (the flop) are revealed, leading to another betting round. This is followed by dealing a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds. The objective is to construct the best five-card poker hand from any combination of hole cards and community cards. Texas Hold’em offers two betting formats: limited betting and no-limit betting. The latter one is noted for its complexity, allowing players to bet any amount up to their entire stack of chips. While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis, we will focus on the algorithms designed to compete in full Texas Hold’em.
@@
Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [10]. It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ [122] to address the issues of storage and computation, respectively, resulting in superhuman performance. Deep CFR [14] combines regret-minimization-based self-play method CFR with deep neural networks. Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method. NSFP [51] introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE. Poker-CNN [140] utilizes self-play to learn convolutional networks to solve a video version of HULHE.
After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ [63]. Thus, traversing the game tree as Cepheus does in HULHE is impossible. DeepStack [95] is a regret-minimization-based self-play method combined with continual re-solving. This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree. Libratus [18] develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR [72]. Recursive Belief-based Learning (ReBeL) [12] introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space. ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks. Unlike the algorithms mentioned above, AlphaHoldem [149] does not use regret-minimization-based self-play. Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings [36], to generate samples.
@@
Pluribus [20], based on Libratus, addresses six-player no-limit Texas Hold’em. Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR [72] to develop its blueprint strategy. It conducts a depth-limited search before executing actions. Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
4.2.2 DouDiZhu